Determine mass spectrometry of following compounds. How do they compare to one another?

1) A compound has the chemical formula C8H15Br. The M+ peak should be at m/z = ____ and there should also be an M+1 peak at m/z = ____ and an M+2 peak at m/z =____ . If the height of the M+ peak is 50, the height of the M+1 peak should be ____ and the height of the M+2 peak should be approximately ____. A compound has the chemical formula C5H11Cl. The M+ peak should be at m/z = ____ and there should also be an M+1 peak at m/z = ____ and an M+2 peak at m/z = ____. If the height of the M+ peak is 27, the height of the M+1 peak should be ____ and the height of the M+2 peak should be approximately____ .
Answer #1

Question: A compound has a chemical formula C8H15Br.

The M+ ion should be the at m/z =   and there should also be an M + 1 peak at m/z=    and an M + 2 peak at m/z=   .

If the height of M+ peak is 50, the height of the M + 1 should be     and the height of the M + 2 should be approximately .

Solution:

Sometion : Ginen chemical formula = Ca H15 Br MW Molecular weight of compound = [no. of atoms x atomic weight of atom). [ccex50 % If af Height of nut is Mtl and M+2 should =? M+1 peak is due to be isotope. 13c is Hly. abundance of total carbons ch co

> Height of M+1 Peak should be 50% of 8.8 = 4.4

Ravi Prabhat Thu, Dec 16, 2021 1:51 AM